# Aharonov-Bohm oscillations and perfectly transmitted mode in amorphous topological insulator nanowires

### v1

@@ IV Fully amorphous nanowire
However, there are additional symmetries which can protect a perfectly transmitted mode, even in fully amorphous nanowires. One possibility is chiral symmetry, which can be present at sufficiently low energies at the surface of the nanowires Xypakis et al. (2020), provided the charge-neutrality point of the surface state is not buried within the bulk bands. Particularly in our model, chiral symmetry is not only present at the surface but it is an exact bulk symmetry, $Sh_{\boldsymbol{r}\boldsymbol{r}^{\prime}}S=-h_{\boldsymbol{r}\boldsymbol{r}^{\prime}}$ with $S=\sigma_{3}\tau_{3}$, at zero Fermi energy and in the absence of scalar disorder. In this limit, the nanowires are in symmetry class AIII Altland and Zirnbauer (1997) and chiral symmetry protects the perfectly transmitted mode.
Even in the absence of chiral symmetry, the perfectly transmitted mode may still be protected in an average sense. Strictly speaking, fully amorphous nanowires break time-reversal symmetry, but for a long enough nanowire the local fluctuations of the flux average out if the amorphicity is spatially isotropic. This situation is realistic for amorphous nanowires since we expect disorder to restore inversion on averageCorbae et al. (2023b); Ciocys et al. (2024); Spring et al. (2021, 2023). Such average symmetry parallels those protecting statistical topological insulators introduced in Ref. Fulga et al., 2014, where it was shown that the surface state remains gapless when subjected to a time-reversal-symmetry-breaking perturbation with zero average, and those protecting against localization in weak topological insulatorsMong et al. (2012); Ringel et al. (2012). Aside from this statistical argument, the effective time-reversal symmetry at the surface is recovered for a fully amorphous nanowire with a large enough cross section. This follows from the fluctuations of the flux $\delta\phi$ approaching zero for increasing linear size $N$ of the cross section, as the fluctuations of the amorphous cross section area scale with the perimeter $\delta\mathcal{A}\sim N$ while the area scales as $\mathcal{A}\sim N^{2}$. Hence, $\delta A/A=\delta\phi/\phi\longrightarrow 0$ as $N$ grows.
Figure 4: (a) Single-realization conductance of a fully amorphous nanowire at $E_{F}=0$ as a function of flux and amorphicity. The parameters are $L=200$, $N_{x}=N_{y}=10$, and $\mu=-1$. (b) Normalized total density of states as a function of the cutoff parameter $R$ for a perfectly transmitted mode (circle), localized mode (triangle) and resonant mode (star) in (a). The meaning of the cutoff $R$ is sketched below the central panel in (c), where $R=0$ corresponds to considering only the central point in the cross section of the nanowires. (c) Local density of states for the perfectly transmitted mode, localized mode and the resonant mode shown in (b).
@@
Fig. 4(b) illustrates the density of states for three representative cases: a perfectly transmitted mode, a localized mode and a resonant mode. These examples are taken from the data in Fig. 4(a), respectively marked as a circle, triangle and star. We take regions delimited by $0<z<L$ and $|x-x^{0}|,|y-y^{0}|<R$, with $(x^{0},y^{0})$ the cross section center, and increase $R$ gradually until it covers the entire cross section of the nanowire. Note that $R=0$ corresponds to the center of the bulk, while $R=5$ corresponds to covering the entire cross section, as $N_{x}=N_{y}=10$. For all three cases only around $10-30\%$ of the density of states is contained within $R\lesssim 3.5$, indicating that transport takes place mostly at the surface. The corresponding local density of states is studied in Fig. 4(c). The perfectly transmitted mode fully hybridizes with both leads without decaying, as backscattering is forbidden by the chiral symmetry. There is an increased local density of states at the hinges of the nanowire, which is expected for topological states Ezawa (2018) and it is induced because the sites in these regions have a lower coordination number than in other parts of the surface. The localized mode decays through the nanowire not hybridizing with the upper lead, which causes a decrease in conductance. Finally, the resonant state hybridizes with the upper lead more faintly than the perfectly transmitted mode, resulting in $G<e^{2}/h$. The local density of states of the resonant state is inhomogeneous along the nanowire, accumulating in the darker regions, and there is no increase along the hinges, contrasting with the perfectly transmitted mode.
@@
Figure 6: Single-realization conductance of a fully amorphous nanowire as a function of Fermi energy and amorphicity. The parameters are $L=200$, $N_{x}=N_{y}=10$, $\mu=0$. The solid lines indicate the conductance at zero magnetic flux while the dotted lines indicate the conductance at the magnetic flux that maximizes conductance at $E_{F}=0.01$. The inset shows the downturn of the $\phi=0$ conductance curves for different cross section sizes as the Fermi energy reaches the bands where the localization length becomes comparable to the cross section linear size, $N$. The nanowires have $N_{x}=N_{y}=N$ and $L=100$. In the $x$-axis, $E_{F}$ is weighted by $N$ to account for the change in level spacing for different cross section sizes.
@@
The difference between the perfectly transmitted mode at low amorphicity and the conductance resonances at high amorphicity is further illustrated in Fig. 5. We plot the single-realization conductance as a function of flux and nanowire length for $w=0.2$ and $w=0.4$. We use a parent nanowire of length $L=200$ and take cuts of varying length. This allows us to keep track of where the resonances appear, as their position is realization dependent. Fig. 5(a) shows that the conductance of the perfectly transmitted mode stays pinned to $G=e^{2}/h$ regardless of the nanowire length, indicating that the mode is never localized. The broadening of the peak for shorter lengths is due to the evanescent modes of the leads that can tunnel through the nanowire, contributing to the conductance. In contrast to the perfectly transmitted mode, Fig. 5(b) shows that, at high amorphicity, the height of most resonances decreases with the nanowire length. Fig. 5(c) shows an approximate exponential decay of three different resonances in Fig. 5(b), indicating the existence of bound states in the wire.
@@ V Topology of amorphous nanowires
We expect the topology of the amorphous nanowires to be protected by the approximate time-reversal symmetry in the average sense discussed in the previous section. However, to characterize the topology of the low-energy regime where the perfectly transmitted mode stands out, we can restrict to the chiral limit. Since the topological protection comes from time-reversal symmetry, adding a time-reversal symmetric perturbation that breaks chiral symmetry, such as adding weak disorder, will not result in a transition to a trivial phase, only in a change of the symmetry class. Chiral symmetry acts then only as an additional symmetry that makes the computation of a local topological marker more direct.
In the chiral limit, fully amorphous nanowires fall in class AIII, and to characterize their topology we use the local chiral marker of Ref. Hannukainen et al., 2022. Local markers are real-space formulations of topological invariants, usually calculated by computing local expectation values of a polynomial of the one-particle density matrix, position and symmetry operators. Averaging a local marker over a sufficiently large bulk region results in a quantized value that characterizes the topological phase of the state. If instead we consider a full system with open boundary conditions, the average local marker over the system necessarily yields zero, as it can generally be expressed as the trace of a commutator, which vanishes in a finite Hilbert space. In the following, we use the Kernel Polynomial Method Weiße et al. (2006); Varjas et al. (2020) to evaluate the one-particle density matrix of the nanowires and compute the associated chiral local marker. We combine this method with the stochastic trace algorithm Weiße et al. (2006); Varjas et al. (2020) when performing averages over large regions of the nanowire.
@@
Figure 7: Realization- and bulk-averaged chiral marker for fully amorphous nanowire as a function of amorphicity and cross section size $N$, with $L=150$. For each realization the local marker is averaged within a region centered at $x_{i}^{0}=(N_{i}-1)/2$ and spanning $|x_{i}-x_{i}^{0}|<0.2N_{i}$ with $x_{1}=x,x_{2}=y$ and $x_{3}=z$, and over $N_{s}=100$ realizations. The shaded regions indicate the standard deviation of the distribution. We used $5000$ moments to calculate the one-particle density matrix with the Kernel Polynomial Method and $5$ random vectors to evaluate the average local marker with the stochastic trace algorithm.
@@
Figure 8: (a-d) Projection onto the $xy$-plane of all sites within a cut spanned by $47.540.5<z<$ 47.5 of a fully amorphous nanowire of length $L=100$ for different amorphicities. Each projection shows $7$ different realizations simultaneously, and the color map indicates the value of the chiral local marker $\nu(\boldsymbol{r})$ at each site. (e-f) Probability distributions for a site located a distance $r=\sqrt{x^{2}+y^{2}}$ from the cross section center having a marker value of $\nu$. The distributions are extracted from the same data shown in the $xy$-projections. The color code indicates the limiting values of $r$ for which sites contribute to $P(\nu)$. The amorphicity values are (a, e) $w=0.1$, (b, f) $w=0.2$, (c, g) $w=0.3$, (d, h) $w=0.4$. We used $1000$ moments to evaluate the one-particle density matrix with the Kernel Polynomial method.
@@
Fig. 7 shows the realization- and bulk-averaged marker $\overline{\nu}$ as a function of amorphicity and cross section size for fully amorphous nanowires of length $L=150$ and $N_{s}=100$ realizations. We also show the standard deviation of the marker distribution obtained for the different amorphous realizations. For each realization we average the local marker over all sites within a region centered at $x_{i}^{0}=(N_{i}-1)/2$ and spanning $|x_{i}-x_{i}^{0}|<0.2N_{i}$ with $x_{1}=x,x_{2}=y$ and $x_{3}=z$. In the crystalline limit, when $w$ is close to zero, the chiral marker approaches the quantized value $\overline{\nu}=-1$ for increasing cross section size, which is expected since for the parameters chosen the crystalline nanowires are in a topological phase. In this limit different realizations of the amorphous structure are very similar to each other, and the only source of statistical uncertainty is in the stochastic trace evaluation. For $w\gtrsim 0.1$ the marker starts deviating from the quantized value and the standard deviation becomes broader. This indicates a topological transition away from the $\overline{\nu}=-1$ phase driven by amorphicity. The transition results in the average marker approaching a noninteger value $\overline{\nu}\approx-0.15$ for $w\gtrsim 0.6$ and increasing cross section size. Presumably, either the single particle density matrix is not exponentially localized in this regime or the cross section size $N$ is shorter than the localization length, making the topological phase of the nanowires not well defined and resulting in a larger statistical uncertainty than at low $w$. Although local marker calculations cannot give more information about this phase at high amorphicity, the analysis of the conductance resonances in Fig. 5(b,c) suggests that the nanowires flow to a trivial insulating phase with an exponentially decaying conductance with the nanowire length, which supports the idea that the localization length of the single particle density matrix exceeds the size of the cross section.
@@
Complementing the average calculation in Fig. 7, the real space distribution of the local marker also signals a topological transition in the nanowires. In Fig. 8(a-d) we show the projection onto the $xy$-plane of the sites contained in $7$ realizations of a cut spanning several cross sections (in the reference lattice) of a nanowire. The color of each site corresponds to $\nu(\boldsymbol{r})$, with $\boldsymbol{r}$ the position in the $xy$-plane. For each projection, we compute the probability distribution $P(\nu)$ of a site situated a distance $r$ from the cross-section center having a certain $\nu$, Fig. 8(e-h). Each vertical panel in the figure corresponds to a different amorphicity. For $w=0.1$ and $w=0.2$ the real space distribution of the marker is the one expected in a topological phase. The nanowires have a topological core with $\nu(\boldsymbol{r})\approx-1$, while the boundary compensates the bulk contribution with $\nu(\boldsymbol{r})>0$. This is explained more thoroughly by the probability distributions $P(\nu)$, which peak at $\nu=-1$ up to $r\approx 6$ for both $w=0.1$ and $w=0.2$. The difference between the two values of $w$ is the spread of the distributions, which broaden with increasing amorphicity. For $r\gtrsim 6$ the distributions start to drift towards higher marker values, eventually peaking and spreading only over positive values of $\nu$. The drift shows how sites at the boundary compensate for the inner topological core, making clear the distinction between bulk sites and boundary sites. For $w=0.3$ the peaks in $P(\nu)$ are not pinned to $\nu=-1$, and although there is a drift, the distributions for the boundary sites generally peak at $\nu=0$ and spread to both negative and positive values. This signals that $\nu(\boldsymbol{r})$ in the bulk is not as homogeneously quantized as for smaller amorphicity, and hence induces less compensation by the boundary. At $w=0.4$ all $P(\nu)$ but for $r<1$ peak around $\nu=0$. The distributions spread mainly to negative values for small $r$, consistent with $\overline{\nu}$ staying below zero for the range of amorphicities considered in Fig. 7, while they spread both to positive and negative values for larger $r$. Although the positive values are still mainly localized at the surface, the distinction between the bulk and the boundary becomes faint, as negative and positive values of $\nu(\boldsymbol{r})$ become less probable than $\nu=0$, see Fig. 8(d), indicating a topological transition.

### v2

@@ IV Fully amorphous nanowire
However, there are additional symmetries which can protect a perfectly transmitted mode, even in fully amorphous nanowires. One possibility is chiral symmetry, which can be present at sufficiently low energies at the surface of the nanowires [71], provided the charge-neutrality point of the surface state is not buried within the bulk bands. Particularly in our model, chiral symmetry is not only present at the surface but it is an exact bulk symmetry, $Sh_{\boldsymbol{r}\boldsymbol{r}^{\prime}}S=-h_{\boldsymbol{r}\boldsymbol{r}^{\prime}}$ with $S=\sigma_{3}\tau_{3}$, at zero Fermi energy and in the absence of scalar disorder. In this limit, the nanowires are in symmetry class AIII [85] and chiral symmetry protects the perfectly transmitted mode [93].
Figure 4: (a) Single-realization conductance of a fully amorphous nanowire at $E_{F}=0$ as a function of flux and amorphicity. The parameters are $L=200$, $N_{x}=N_{y}=10$, and $\mu=-1$. (b) Total density of states per unit area as a function of the cutoff parameter $R$ for a perfectly transmitted mode (circle), localized mode (triangle) and resonant mode (star) in (a). The meaning of the cutoff $R$ is sketched below the central panel in (c), where $R=0$ corresponds to considering only the central point in the cross section of the nanowires. Each curve is normalized to its total density of states per area at $R=5$. (c) Local density of states for the perfectly transmitted mode, localized mode and the resonant mode shown in (b).
Even in the absence of chiral symmetry, the perfectly transmitted mode may still be protected in an average sense. Strictly speaking, fully amorphous nanowires break time-reversal symmetry, but for a long enough nanowire the local fluctuations of the flux average out if the amorphicity is spatially isotropic. This situation is realistic for amorphous nanowires since we expect disorder to restore inversion on average [62; 63; 94; 95]. Such average symmetry parallels those protecting statistical topological insulators introduced in Ref. 96, where it was shown that the surface state remains gapless when subjected to a time-reversal-symmetry-breaking perturbation with zero average, and those protecting against localization in weak topological insulators [97; 98]. Aside from this statistical argument, the effective time-reversal symmetry at the surface is recovered for a fully amorphous nanowire with a large enough cross section. This follows from the fluctuations of the flux $\delta\phi$ approaching zero for increasing linear size $N$ of the cross section, as the fluctuations of the amorphous cross section area scale with the perimeter $\delta\mathcal{A}\sim N$ while the area scales as $\mathcal{A}\sim N^{2}$. Hence, $\delta A/A=\delta\phi/\phi\longrightarrow 0$ as $N$ grows.
@@
Figure 6: Single-realization conductance of a fully amorphous nanowire as a function of Fermi energy and amorphicity. The parameters are $L=200$, $N_{x}=N_{y}=10$, $\mu=0$. The solid lines indicate the conductance at zero magnetic flux while the dotted lines indicate the conductance at the magnetic flux that maximizes conductance at $E_{F}=0.01$. The inset shows the downturn of the $\phi=0$ conductance curves for different cross section sizes as the Fermi energy reaches the bands where the localization length becomes comparable to the cross section linear size, NN. The nanowires have $N_{x}=N_{y}=N$ and $L=100$. In the $x$-axis, $E_{F}$ is weighted by NN to account for the change in level spacing for different cross section sizes.
Fig. 4(b) illustrates the density of states per unit area for three representative cases: a perfectly transmitted mode, a localized mode and a resonant mode. These examples are taken from the data in Fig. 4(a), respectively marked as a circle, triangle and star. We take regions delimited by $0<z<L$ and $|x-x^{0}|,|y-y^{0}|<R$, with $(x^{0},y^{0})$ the cross section center, and increase $R$ gradually until it covers the entire cross section of the nanowire. Note that $R=0$ corresponds to the center of the bulk, while $R=5$ corresponds to covering the entire cross section, as $N_{x}=N_{y}=10$. The density of states in the perfectly transmitted and localized modes is mostly located in the surface, while in the resonant mode the penetration length into the bulk becomes larger due to the strong amorphicity. The corresponding local density of states is studied in Fig. 4(c). The perfectly transmitted mode fully hybridizes with both leads without decaying, as backscattering is forbidden by the chiral symmetry. There is an increased local density of states at the hinges of the nanowire, which is expected for topological states [100] and it is induced because the sites in these regions have a lower coordination number than in other parts of the surface. The localized mode decays through the nanowire not hybridizing with the upper lead, which causes a decrease in conductance. Finally, the resonant state hybridizes with the upper lead more faintly than the perfectly transmitted mode, resulting in $G<e^{2}/h$. The local density of states of the resonant state is inhomogeneous along the nanowire, accumulating in the darker regions, and there is no increase along the hinges, contrasting with the perfectly transmitted mode.
@@
The difference between the perfectly transmitted mode at low amorphicity and the conductance resonances at high amorphicity is further illustrated in Fig. 5. We plot the single-realization conductance as a function of flux and nanowire length for $w=0.2$ and $w=0.4$. We use a parent nanowire of length $L=200$ and take cuts of varying length. This allows us to keep track of where the resonances appear, as their position is realization dependent. Fig. 5(a) shows that the conductance of the perfectly transmitted mode stays pinned to $G=e^{2}/h$ regardless of the nanowire length, indicating that the mode is never localized. The broadening of the peak for shorter lengths is due to the evanescent modes of the leads that tunnel through the nanowire with a decay length that scales as $\ell_{\rm tunnel}\sim N/L$ [101], contributing to the conductance. In contrast to the perfectly transmitted mode, Fig. 5(b) shows that, at high amorphicity, the height of most resonances decreases with the nanowire length. Multiple evanescent modes may contribute to the conductance in this regime, especially for short nanowires, where the conductance of certain resonances becomes $G>e^{2}/h$. Fig. 5(c) shows an approximate exponential decay of three different resonances in Fig. 5(b), indicating the existence of bound states in the wire.
@@ V Topology of amorphous nanowires
We expect the topology of the amorphous nanowires to be protected by the approximate time-reversal symmetry in the average sense discussed in the previous section. However, to characterize the topology of the low-energy regime where the perfectly transmitted mode stands out, we can restrict to the chiral limit. Since the topological protection comes from time-reversal symmetry, adding a time-reversal symmetric perturbation that breaks chiral symmetry, such as adding weak disorder, will not result in a transition to a trivial phase, only in a change of the symmetry class. Chiral symmetry acts then only as an additional symmetry that facilitates the topological characterization of the nanowires, as it makes the computation of a local topological marker more direct.
@@
Figure 7: Realization- and bulk-averaged chiral marker for fully amorphous nanowire as a function of amorphicity and cross section size $N$, with $L=150$. For each realization the local marker is averaged within a region centered at $x_{i}^{0}=(N_{i}-1)/2$ and spanning $|x_{i}-x_{i}^{0}|<0.2N_{i}$ with $x_{1}=x,x_{2}=y$ and $x_{3}=z$, and over $N_{s}=100$ realizations. The shaded regions indicate the standard deviation of the distribution. We used 50005000 moments to calculate the one-particle density matrix with the Kernel Polynomial Method and 55 random vectors to evaluate the average local marker with the stochastic trace algorithm. The inset shows an analogous calculation where we set $\epsilon=3$ (in the main panel $\epsilon=4$).
@@
Figure 8: (a-d) Projection onto the $xy$-plane of all sites within a cut spanned by $40.5<z<47.5$ of a fully amorphous nanowire of length $L=100$ for different amorphicities. Each projection shows 77 different realizations simultaneously, and the color map indicates the value of the chiral local marker $\nu(\boldsymbol{r})$ at each site. (e-f) Probability distributions for a site located within an inner and an outer concentric square in the $xy$-plane having a marker value of $\nu$. The parameter $R$ controls the size of the inner and outer squares as in Fig. 4, and the color code indicates within which concentric squares sites contribute to $P(\nu)$. The distributions are extracted from the same data shown in the xyxy-projections. The amorphicity values are (a, e) $w=0.1$, (b, f) $w=0.2$, (c, g) $w=0.3$, (d, h) $w=0.4$. We used 10001000 moments to evaluate the one-particle density matrix with the Kernel Polynomial method.
In the chiral limit, fully amorphous nanowires fall in class AIII, and to characterize their topology we use the local chiral marker of Ref. 80, given by
[…]
Local markers are real-space formulations of topological invariants, usually calculated by computing local expectation values of a polynomial of the one-particle density matrix, position and symmetry operators, as in Eq. 6. Averaging a local marker over a sufficiently large bulk region results in a quantized value that characterizes the topological phase of the state. If instead we consider a full system with open boundary conditions, the average local marker over the system necessarily yields zero, as it can generally be expressed as the trace of a commutator, which vanishes in a finite Hilbert space. In the following, we use the Kernel Polynomial Method [102; 103] to evaluate the one-particle density matrix of the nanowires and compute the associated chiral local marker. We combine this method with the stochastic trace algorithm [102; 103] when performing averages over large regions of the nanowire.
@@
Fig. 7 shows the realization- and bulk-averaged marker $\overline{\nu}$ as a function of amorphicity and cross section size for fully amorphous nanowires of length $L=150$ and $N_{s}=100$ realizations. We also show the standard deviation of the marker distribution obtained for the different amorphous realizations. For each realization we average the local marker over all sites within a region centered at $x_{i}^{0}=(N_{i}-1)/2$ and spanning $|x_{i}-x_{i}^{0}|<0.2N_{i}$ with $x_{1}=x,x_{2}=y$ and $x_{3}=z$. In the crystalline limit, when $w$ is close to zero, the chiral marker approaches the quantized value $\overline{\nu}=-1$ for increasing cross section size, which is expected since for the parameters chosen the crystalline nanowires are in a topological phase. In this limit different realizations of the amorphous structure are very similar to each other, and the only source of statistical uncertainty is in the stochastic trace evaluation. For $w\gtrsim 0.1$ the marker starts deviating from the quantized value and the standard deviation becomes broader. This indicates a topological transition away from the $\overline{\nu}=-1$ phase driven by amorphicity, similar to the disorder- and impurity-driven transitions studied in crystalline structures [104; 105]. The transition results in the average marker approaching a noninteger value $\overline{\nu}\approx-0.15$ for $w\gtrsim 0.6$ and increasing cross section size. Presumably, either the single particle density matrix is not exponentially localized in this regime or the cross section size $N$ is shorter than the localization length, making the topological phase of the nanowires not well defined and resulting in a larger statistical uncertainty than at low $w$. We stress that this situation is not an artifact of our calculation and holds physical meaning: due to their finite cross section, nanowires may not have a well defined topological phase even if the latter is well defined in the thermodynamic limit. These finite size effects are in-built and shape the physics of the nanowires, rather than only being of numerical importance. We have further checked (data not shown) that the approach to a non-well-defined topological phase does not change by fixing the aspect ratio of the nanowires, ruling out the possibility of a dimensional crossover effect obscuring the results. Although local marker calculations cannot give more information about this phase at high amorphicity, the analysis of the conductance resonances in Fig. 5(b,c) suggests that the nanowires flow to a trivial insulating phase with an exponentially decaying conductance with the nanowire length, which supports the idea that the localization length of the single particle density matrix exceeds the size of the cross section. Whether the amorphicity-driven transition converges to a well-defined trivial phase depends on the exact parameters of the calculation. We exemplify this in the inset of Fig. 7, where changing the parameter $\epsilon$ results in a topological transition that converges to a trivial phase $\overline{\nu}=0$ at large $w$. In this particular case, the bulk band gap in the crystalline limit is reduced as compared to the main panel in Fig. 7, which may explain why the range of amorphicities considered is enough to trivialize the nanowires as opposed to the main panel.
@@
Complementing the average calculation in Fig. 7, the real space distribution of the local marker also signals a topological transition in the nanowires. In Fig. 8(a-d) we show the projection onto the $xy$-plane of the sites contained in $7$ realizations of a cut spanning several cross sections (in the reference lattice) of a nanowire. The color of each site corresponds to $\nu(\boldsymbol{r})$, with $\boldsymbol{r}$ the position in the $xy$-plane. For each projection, we compute the probability distribution $P(\nu)$ of a site situated within an inner and an outer concentric square having a certain $\nu$. We vary the sizes of the squares, controlled by the $R$ parameter defined in Fig. 4, in order to obtain different probability distributions associated with different regions in the $xy$-plane, from regions including sites very close to the center of the cross section to regions including sites only around the boundary of the nanowires. The resulting probability distributions are shown in Fig. 8(e-h). Each vertical panel in the figure corresponds to a different amorphicity. For $w=0.1$ and $w=0.2$ the real space distribution of the marker is the one expected in a topological phase. The nanowires have a topological core with $\nu(\boldsymbol{r})\approx-1$, while the boundary compensates the bulk contribution with $\nu(\boldsymbol{r})>0$. This is explained more thoroughly by the probability distributions $P(\nu)$, which peak at $\nu=-1$ up to $R\simeq 5$ for both $w=0.1$ and $w=0.2$. Although peaked at $\nu=-1$, the probability distributions for small $R$ broaden with increasing amorphicity. At $R\gtrsim 5$ for $w=0.1$ and $w=0.2$ the distributions start to drift towards higher marker values, eventually peaking and spreading over positive values of $\nu$. The drift shows how sites at the boundary compensate for the inner topological core, making clear the distinction between bulk sites and boundary sites. For $w=0.3$ the peaks in $P(\nu)$ are not pinned to $\nu=-1$, and although there is a drift, the distributions for the boundary sites generally peak at $\nu=0$ and spread to both negative and positive values. This signals that $\nu(\boldsymbol{r})$ in the bulk is not as homogeneously quantized as for smaller amorphicity, and hence induces less compensation by the boundary. At $w=0.4$ all $P(\nu)$ peak around $\nu=0$. The distributions spread mainly to negative values for small $R$, consistent with $\overline{\nu}$ staying below zero for the range of amorphicities considered in Fig. 7, while they spread both to positive and negative values for larger $R$. Although the positive values are generally localized close to the surface, the distinction between the bulk and the boundary becomes faint, as negative and positive values of $\nu(\boldsymbol{r})$ become less probable than $\nu=0$, see Fig. 8(d), indicating a topological transition.